Two cars travel in opposite directions from the same starting point. The rate of one car is 20 miles per hour less that the other car. After 5 hours the cars were 700 mile apart. what was the rate of each car?
Answer by checkley77(12844)

5x+5(x-20)=700
5x+5x-100=700
10x=700+100
10x=800
x=800/10
x=80 mph for the faster car.
80-20=60 mph for the slower car.
Proof:
5*80+5*60=700
400+300=700
700=700